OpenStudy (anonymous):

100 is 250% of ___a0. plezz help :(

OpenStudy (anonymous):

(250/100) x = 100 so x = 100/(250/100) x = 40

OpenStudy (anonymous):

wait no its not?

OpenStudy (anonymous):

what do you mean by that?